Direct Labor Variances

Bellingham Company produces a product that requires 2 standard direct labor hours per unit at a standard hourly rate of $18.00 per hour. If 3,600 units used 7,400 hours at an hourly rate of $17.46 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $
b. Direct labor time variance $
c. Direct labor cost variance $
Answer #1
a. Direct labor rate variance -3996 Favorable =7400*(17.46-18)
b. Direct labor time variance 3600 Unfavorable =18*(7400-3600*2)
c. Direct labor cost variance -396 Favorable =(7400*17.46)-(3600*2*18)